OpenStudy (anonymous):

Solve by substitution method: y = -2x + 3 3x - 2y = 8 A. no solution B. (14, -25) C. (-14, 31) D. (2, -1)

jhonyy9 (jhonyy9):

hi ! do you know what mean substitution ?

jhonyy9 (jhonyy9):

so in the first equation there is a value of y than just need to substitute it in second equation in place of y and calculi the x ok ?

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

so after that you just substitute the value of x inside one equation in place of x and will get the value of y ok ?

jhonyy9 (jhonyy9):

so let you try ,i like see your calculi ok ?

OpenStudy (texaschic101):

y = -2x + 3 now sub -2x + 3 in for y in the other equation 3x - 2y = 8 3x - 2(-2x + 3) = 8 can you take it from here ?

OpenStudy (anonymous):

x = 2

OpenStudy (texaschic101):

3x - 2(-2x + 3) = 8 3x + 4x - 6 = 8 3x + 4x = 8 + 6 7x = 14 x = 2 correct. Now sub 2 in for x in either of the equations to find y

OpenStudy (anonymous):

y = -2x + 3 3x - 2y = 8 3x - 2 (-2x+3)=8 3x +4x-6=8 7x-6=8/+6 7x=14/:7 x=2 y=-2x+3 y=-2 * 2 +3 y=-4 + 3 y=-1

OpenStudy (texaschic101):

oh well..nvm...you just got the answer handed to you. I hope you fully understand how to do these types of problems.